# Neural network learns low-dimensional polynomials with SGD near the information-theoretic limit

### v1

@@ B.3.2 When $\mathrm{IE}(\sigma_{*}^{I})=1$ with $I\geq 2$
$\boldsymbol{g}=y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}.$ (B.75)
From (B.37) and (B.38) we know that the first term of $\boldsymbol{g}$ (B.75) is evaluated as
$\mathbb{E}[y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]=\sum_{j=p}^{q}j!\alpha_{j}\beta_{j}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j-1}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{\theta},$ (B.76)
where the sum of coefficients are bounded by $\lesssim\kappa$.
We then consider the second term of $\boldsymbol{g}$ (B.75). We can replace $\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}$ by $d$ with the following bound similarly to (B.40):
$\big{|}\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]-\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]\big{|}\leq C_{1}\eta d^{\frac{1}{2}}.$ (B.77)
For the term $\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$. the following decomposition can be made.
$\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ (B.78)
$=\sum_{j=1}^{C_{\sigma}}j\beta_{j}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}$ (B.79)
$=\sum_{k=0}^{j-1}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose k}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}^{k}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}.$ (B.80)
We evaluate each term of (B.80) except for $k=I-1$. Similarly to the bounds on (B.46), each term is a constant multiple of $\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}$, and we want to bound
$\bigg{|}\boldsymbol{\theta}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}^{k}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}.$ (B.81)
When $k\leq I-2$, $\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}),\cdots,\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})^{k+1}$ has information exponent larger than $1$. Therefore, we have
(B.81)$\lesssim(\eta d)^{k}\kappa.$ (B.82)
When $k\geq I$, we have
(B.81)$\lesssim(\eta d)^{k}\leq c_{1}^{I}.$ (B.83)
Now we consider the case when $k=I-1$:
$\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})+\upsilon\big{)}\big{)}^{I}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}$ (B.84)
$=\sum_{l=0}^{I}\sum_{j=1}^{C_{\sigma}}(\eta d)^{k}j\beta_{j}{j-1\choose I-1}{I\choose l}\mathbb{E}[\upsilon^{l}]\mathbb{E}\bigg{[}\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{I-1}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{I-l}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}.$ (B.85)
Note that $\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}),\cdots,(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{I-1}$ has information exponent larger than $1$. Therefore, for $l\geq 1$, we have
$\bigg{|}\boldsymbol{\theta}^{\top}(\eta d)^{I-1}j\beta_{j}{j-1\choose I-1}{I\choose l}\mathbb{E}[\upsilon^{l}]\mathbb{E}\bigg{[}\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{I-1}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{I-l}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}$ (B.86)
$\lesssim(\eta d)^{I-1}\bigg{|}\mathbb{E}\bigg{[}\sum_{m=2}^{\infty}m!\kappa^{m-1}H\big{(}\mathsf{He}_{j-I}(\sigma^{\prime})^{I-1};m-1\big{)}H\big{(}\big{(}\sigma_{*}\big{)}^{I-l};m\big{)}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}$ (B.87)
$\lesssim(\eta d)^{I-1}\kappa.$ (B.88)
And for $l=0$, $(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{I}$ has information exponent $1$ and we have
$\sum_{j=1}^{C_{\sigma}}(\eta d)^{I-1}j\beta_{j}{j-1\choose I-1}\mathbb{E}\bigg{[}\mathsf{He}_{j-I}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{I-1}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{I-l}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}$ (B.89)
$=(\eta d)^{I-1}\sum_{m=1}^{\infty}m!\kappa^{m-1}H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};m-1\bigg{)}H\big{(}(\sigma_{*})^{I};m\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}$ (B.90)
If $H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};0\bigg{)}\neq 0$, we have
(B.90)$=(1+O(\kappa))\cdot(\eta d)^{I-1}\underbrace{H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};0\bigg{)}H\big{(}(\sigma_{*})^{I};1\big{)}}_{=:\gamma}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}.$ (B.91)
Note that
$H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};0\bigg{)}=H\bigg{(}\frac{1}{(I-1)!}\sigma^{(I)}(\sigma^{(1)})^{I-1};0\bigg{)}.$ (B.92)
Now we have that
$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ (B.93)
$=(1+O(\kappa))\cdot(\eta d)^{I-1}\gamma\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+O(\kappa+C_{1}\eta d^{\frac{1}{2}}+(\eta d)^{I-2}\kappa+(\eta d)^{I}+(\eta d)^{I-1}\kappa).$ (B.94)
When $d^{-\frac{1}{2}}C_{3}\leq\kappa\leq c_{2}$ and $\eta=c_{1}d^{-1}$, we have
$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]=(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma$ (B.95)
Together with the bound on the first term ((B.76)), the expectation of $\boldsymbol{g}$ is evaluated as
$\boldsymbol{\theta}^{\top}\mathbb{E}[\boldsymbol{g}]=(1+O(c_{1}))\cdot c_{1}^{I-1}.$ (B.96)
By using this and (B.73), the expected update of the alignment becomes
$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]$ $\geq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma-\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\mathbb{E}[\|\boldsymbol{g}\|^{2}]-\frac{1}{2}\eta^{3}\tilde{\eta}^{3}\mathbb{E}[|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}].$ (B.97)
Note that $\mathbb{E}[\|\boldsymbol{g}\|^{2}],\mathbb{E}[|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}]\lesssim d$ and $\kappa\lesssim c_{1}$. Taking $\tilde{\eta}\leq c_{2}$ yields that
$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]$ $\geq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma.$ (B.98)
The upper bound can be obtained in similar fashion,
$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]$ $\leq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma.$ (B.99)
Finally, the noise can be handled in the exact same way as that of Lemma 9, the details of which we omit. ∎
B.3.3 When $\mathrm{IE}(\sigma_{*}^{I})=1,2$ with $I=1$
We finally consider the case when $\mathrm{IE}[\sigma_{*}]=1$ or when $\mathrm{IE}[\sigma_{*}]=2$ and $\mathrm{IE}[\sigma_{*}^{i}]\geq 2$ for $i=2,\cdots,C_{\sigma}$.

### v2

@@ B.5 From Weak Recovery to Strong Recovery
$\boldsymbol{g}=P_{\boldsymbol{w}^{2t}}y\sigma^{\prime}({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{x})\boldsymbol{x}.$ (B.109)
From (B.106), the expectation of (B.107) is bounded by
$\mathbb{E}[\kappa^{2(t+1)}]$ $\geq\kappa^{2t}+\eta(1-(\kappa^{2t})^{2})\sum_{j=p}^{\infty}j!\alpha_{j}\beta_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t})^{j-1}-\eta^{2}C_{4}d(\kappa^{2t}+\eta)$ (B.110)
$\geq\kappa^{2t}+\eta(1-(\kappa^{2t})^{2})\sum_{j=p}^{\infty}j!\alpha_{j}\beta_{j}(\kappa^{2t})^{p-1}-\eta^{2}C_{4}d(\kappa^{2t}+\eta).$ (B.111)
By letting $\eta\leq c_{1}^{p-1}\varepsilon d^{-1}$, when $\kappa^{2t}\leq 1-\varepsilon$, we have
$\mathbb{E}[\kappa^{2(t+1)}]\geq\kappa^{2t}+\frac{1}{2}\eta\varepsilon\sum_{j=p}^{\infty}j!\alpha_{j}\beta_{j}(\kappa^{2t})^{p-1}\geq\kappa^{2t}+\eta\varepsilon c_{1}^{p}.$ (B.112)
It is easy to see that the noise $\nu^{2t}$ has sub-Weibull tail and we obtain that
$\kappa^{2(t+1)}\geq\kappa^{2t}+\frac{1}{2}\eta\varepsilon\sum_{j=p}^{\infty}j!\alpha_{j}\beta_{j}(\kappa^{2t})^{p-1}+\eta\nu^{2t}\geq\kappa^{2t}+\eta\varepsilon c_{1}^{p}+\eta\nu^{2t}.$ (B.113)
Suppose that $2c_{2}\leq\kappa^{2(T{1,1}+\tau)}\leq 1-\varepsilon$ for all $t=0,1,\dots,\tau-1$. By taking the summation of (B.113), we have
$\kappa^{2(T{1,1}+\tau)}\geq\kappa^{2T{1,1}}+\eta\varepsilon tc_{1}^{p}+\eta\sum_{s=T{1,1}}^{T{1,1}+\tau-1}\nu^{2t}\geq 2c_{2}+\eta\varepsilon\tau c_{1}^{p}-C_{4}\eta\sqrt{\tau},$ (B.114)
with high probability. The third term is bounded by $C_{4}\eta\sqrt{\tau}\leq c_{2}$ when $\tau\leq c_{2}^{2}C_{4}^{-2}\eta^{-2}=c_{2}^{2}C_{4}^{-2}\bar{c}_{\eta}^{-2}\varepsilon^{-2}d^{2}$ and by $\frac{1}{2}\eta\varepsilon\tau c_{1}^{p}$ when $\tau\geq 4\varepsilon^{-2}c_{1}^{-2p}C_{4}^{2}$. Because $c_{2}^{2}C_{4}^{-2}\bar{c}_{\eta}^{-2}\varepsilon^{-2}d^{2}\geq 4\varepsilon^{-2}c_{1}^{-2p}C_{4}^{2}$, we can bound (B.114) by
$\kappa^{2(T{1,1}+\tau)}\geq c_{2}+\frac{1}{2}\eta\varepsilon\tau c_{1}^{p},$ (B.115)
which verifies $2c_{2}\leq\kappa^{2(T{1,1}+\tau)}$.
Therefore, by induction, until $\kappa^{2t}\geq 1-\varepsilon$, we have the lower bound (B.115), whose RHS exceeds $1-\varepsilon$ when $\tau\geq 2\eta^{-1}\varepsilon^{-1}c_{1}^{-p}\leq C_{3}d\varepsilon^{-2}$. Thus, there exists $\tau_{*}\leq T_{1,2}=C_{3}d\varepsilon^{-2}$ such that $\kappa^{2(T{1,1}+\tau_{*})}\geq 1-\varepsilon$, with high probability.
Now, what remains is to prove that $\kappa^{2(T{1,1}+\tau)}\geq 1-3\varepsilon$ holds for all $\tau_{*}\leq t\leq T_{1,2}=C_{3}d\varepsilon^{-2}$. Because $\nu^{2t}$ are mean-zero sub-Weibull random variables, we have that $|\sum_{s=\tau}^{\tau+\tau^{\prime}-1}\nu^{2s}|\leq C_{4}\sqrt{\tau^{\prime}}$ for all $0\leq\tau,\tau^{\prime}\leq T_{1,1}$ with high probability. Also, because $\eta^{t}\ll\varepsilon d^{-1}$, we can easily see that $|\kappa^{2(\tau+1)}-\kappa^{2\tau}|=\tilde{O}(\varepsilon d^{-1})$ for all $\tau=0,1,\cdots,T_{1,1}-1$, with high probability. Thus, when there exists $\tau\geq\tau_{*}$ such that $\kappa^{2(T_{1,1}+\tau-1)}\geq 1-\varepsilon$ and $\kappa^{2(T_{1,1}+\tau)}<1-\varepsilon$, we have $\kappa^{2(T_{1,1}+\tau)}\geq 1-2\varepsilon$ with high probability. Moreover, following the above argument, we can inductively show that
$\kappa^{2(T{1,1}+\tau+\tau^{\prime})}$ $\geq 1-2\varepsilon+\eta\varepsilon\tau^{\prime}c_{1}^{p}-C_{4}\eta\sqrt{\tau^{\prime}}$ (B.116)
$\geq 1-2\varepsilon+\eta\varepsilon\tau^{\prime}c_{1}^{p}-\begin{cases}\varepsilon&(\tau^{\prime}\leq\bar{c}_{\eta}^{-2}C_{4}^{-2}d^{2})\\ \eta\varepsilon\tau^{\prime}c_{1}^{p}&(\tau^{\prime}\geq\varepsilon^{-2}C_{4}^{2}c_{1}^{-2p})\end{cases}.$ (B.117)
$\geq 1-3\varepsilon,$ (B.118)
for $\tau^{\prime}\leq T_{1,2}$ or until $\kappa^{2(T{1,1}+\tau+\tau^{\prime})}\geq 1-\varepsilon$ holds again. Note that the last inequality follows from $\bar{c}_{\eta}^{-2}C_{4}^{-2}d^{2}\geq\varepsilon^{-2}C_{4}^{2}c_{1}^{-2p}$. By repeating this argument (if there are multiple such $\tau$), we can see that $\kappa^{2(T{1,1}+t)}\geq 1-\varepsilon$ holds for all $\tau_{*}\leq t\leq T_{1,2}=C_{3}d\varepsilon^{-2}$ with high probability.
Adjusting hidden constants to remove a factor of $3$ from $3\varepsilon$ yields the desired result. ∎
B.6 Second Layer Training
From the previous analysis, we know that at least $\Omega(1)$ portion of the neurons will satisfy the weak and strong recovery conditions (Appendix B.1), at least $\tilde{\Omega}(1)$ portion of the neurons (independent from the choice of $\sigma_{j}$) satisfy initial alignment conditions (Appendix B.2), and at least $1-o(1)$ fraction of them achieves strong recovery. This subsection proves a generalization error bound after second-layer training. Let $f_{\boldsymbol{a}}(\boldsymbol{x})=f_{\boldsymbol{\Theta}}(\boldsymbol{x})$ for $\boldsymbol{\Theta}=(\hat{\boldsymbol{w}}_{j},a_{j},\hat{b}_{j})_{j=1}^{N}$ where $\boldsymbol{a}\in\mathbb{R}^{N}$ and $(\hat{\boldsymbol{w}}_{j},\hat{b}_{j})_{j=1}^{N}$ are the parameters trained in the first stage. Let $\boldsymbol{a}^{*}\in\mathbb{R}^{N}$ be the “certificate” with $\|\boldsymbol{a}^{*}\|^{2}=\tilde{O}(N)$ that is shown to exist in Lemma 22.
Polynomial Link Functions.
The following lemma is a complete version of Proposition 5.